### v1

@@ Abstract
This paper introduces PublicHearingBR, a Brazilian Portuguese dataset designed for summarizing long documents. The dataset consists of transcripts of public hearings held by the Brazilian Chamber of Deputies, paired with news articles and structured summaries containing the individuals participating in the hearing and their statements or opinions. The dataset supports the development and evaluation of long document summarization systems in Portuguese. Our contributions include the dataset, a hybrid summarization system to establish a baseline for future studies, and a discussion on evaluation metrics for summarization involving large language models, addressing the challenge of hallucination in the generated summaries. As a result of this discussion, the dataset also provides annotated data that can be used in Natural Language Inference tasks in Portuguese.
@@ 1 Introduction
With huge volumes of text being added daily to information systems, it is difficult and time-consuming for users to search through that content [3]. Textual data often contains irrelevant information mixed with relevant information [1]. Thus, in many contexts, it is easier to access a summary containing only the most relevant information of the data.
@@
Text summarization is a task in natural language processing that condenses text into a shorter version while retaining its main ideas. This task is important for various applications, including information retrieval, content management, and automatic news article writing based on documents, among others. Manual summarization is costly and, in many cases, impractical. Therefore, automatic text summarization is necessary [3].
@@
Automatic text summarization methods are categorized into three types: abstractive, extractive, and hybrid [11, 10]. Abstractive summarization generates new sentences that capture the essence of the original text. Extractive summarization selects and concatenates segments directly from the source text. Hybrid method combines elements of both methods. Traditionally, most research has focused on extractive methods. However, in recent years, the emphasis has shifted to abstractive or hybrid methods [3]. With advancements in generative artificial intelligence systems, these methods have become even more significant.
@@
The type of documents involved in summarization tasks can vary, ranging from short news articles to long research papers and collections of multiple documents. Each type presents distinct challenges. Summarizing short documents (such as a news article) is easier to perform and evaluate. Summarizing multiple documents typically depends on a prior retrieval stage on a specific topic within those documents. Summarizing long documents usually involves high compression rates, which means it is necessary to assess what the most relevant pieces of information in the document are. As the information may be dispersed throughout the text, it can be a complex task.
@@
A typical pipeline for evaluating an automatic text summarization system includes some stages: pre-processing, summarization, post-processing, and evaluation [10]. The system evaluation is done by comparing the generated summary with a model summary, using metrics such as ROUGE [7] and Check-Eval [14]. Regardless of the evaluation metric used, public standardized datasets are generally used for the evaluation of new methods.
@@
The availability of datasets is fundamental to advance research and development in summarization methods [5], as they provide the necessary resources to train and evaluate new systems. However, most available datasets are in English. This limitation hinders the development of summarization tools for non-English speakers. Portuguese, one of the world’s most spoken languages [^1] https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers, exemplifies this fact. Despite its wide usage, there is a scarcity of summarization datasets in Portuguese, and existing datasets usually focus on short documents like news articles, leaving a gap in resources for long document summarization.
@@
To address this gap, this paper introduces PublicHearingBR, an open dataset designed for the summarization of long documents in Brazilian Portuguese. The documents to be summarized are transcripts of public hearings held by the Brazilian Chamber of Deputies. A public hearing is an event that can last several hours, during which individuals are called to discuss a specific topic. The transcript of each hearing is published online. Among the characteristics of these documents that may pose challenges for summarization, we list the following: 1. The transcripts are extensive texts, containing dozens of pages and sometimes exceeding a hundred; 2. The transcripts contain the opinions of some individuals on the topic under discussion, meaning that what one person says may be dispersed throughout the transcript; 3. There may also be agreement or disagreement among participants, implying that some opinions of a participant may be inferred from the agreement or disagreement with other participants; and 4. The topics discussed are diverse, ranging from socially controversial issues (e.g., vaccination mandates, minority rights, and climate issues) to highly specific topics (e.g., regulation of online platforms and installation of offshore wind farms).
@@
PublicHearingBR pairs the transcripts with corresponding news articles and metadata indicating the individuals participating in the hearing and their opinions (a structured summary). The articles were extracted from the Agência Câmara de Notícias [^2] https://www.camara.leg.br/noticias, the news agency of the Brazilian Chamber of Deputies. Thus, we consider their text to be adequate summaries of the hearings. The metadata was extracted from the articles.
@@
To test the dataset, we created a custom GPT [^3] https://openai.com/index/introducing-gpts/ to extract relevant information from the transcripts using a hybrid summarization approach. The goal is not to solve the problem of summarizing public hearings, but rather to establish a baseline using a tool available to end users. To evaluate the result, we propose using large language models (LLMs) to compare the results of the experiment with the dataset. We use LLMs both to verify if the extracted opinions are present in the dataset (similar to the Check-Eval metric [14]) and to determine if they are hallucinations.
The contributions of this paper are: 1. PublicHearingBR, the dataset for the summarization of long documents in Portuguese; 2. a LLM-based hybrid summarization system, which provides a baseline for future studies; 3. a discussion on evaluation metrics for summarization using LLMs, including how to identify hallucinations in the generated summary.
@@
The discussion about hallucinations also resulted in manually annotated data containing 4 238 opinions, chunks of text, and a flag indicating whether the opinion can be inferred from those excerpts, i.e., PublicHearingBR can also be used as a Natural Language Inference (NLI) dataset in Brazilian Portuguese.
@@ 2 Related Work
The availability of datasets in Portuguese is limited. There are large corpora in Brazilian Portuguese that have been published recently (e.g., [15]), but they are suitable for other tasks (e.g., fine-tuning LLMs). There are few datasets for text summarization in Portuguese, and they typically focus on summarizing short texts.
@@
One of the first summarization datasets available in Portuguese is the TeMário [13], which contains 100 pairs of journalistic news extracted from two newspapers and summaries written by a specialist. On average, the texts contain approximately 600 words, and the summaries are about 200 words. When writing the summary, the specialist was explicited told to keep its length as 25% to 30% of the original article. The dataset was later expanded with an additional 150 samples, with 1 200 words per text on average [9].
@@
Other datasets in Portuguese for news summarization are the CSTNews [2] and the RecognaSumm [12]. CSTNews is a manually annotated dataset containing 140 samples of short texts. All the texts together sum to about 47 000 words, that is, just over 330 words per sample. On the other hand, RecognaSumm has about 130 000 samples and considers that the summary is the heading and subheading of the news. However, this information does not necessarily translate into a good summary of a text, as the heading and subheading are usually created to describe the main topic of the news to capture the reader’s attention, rather than to serve as an effective summary of the article.
@@
News articles are, in general, short texts. A Portuguese dataset with slightly larger documents is the RulingBR [16]. This dataset contains more than 10 000 rulings from the Brazilian Supreme Federal Court. Each document contains four sections: Ementa, Relatório, Voto, and Acórdão. The Ementa section is considered the reference summary of the decision. In this dataset, the documents contain an average of 1 200 words, reaching up to 2 000 words. The summaries contain up to 150 words, with an average of about 90 words, making the summary about 7% of the total document size. This dataset focuses on summarizing the document to find what the decision is.
@@
There are also multilingual summarization datasets, notably WikiLingua [6] and XL-Sum [4]. The former deals with summarizing "How To" texts, while the latter focuses on summarizing news articles. Both contain text pairs in English and summaries in one or more languages (up to 18 languages for WikiLingua and up to 44 languages for XL-Sum). Generally, the texts are short. For instance, the average length of WikiLingua articles is 391 tokens (using the mBART tokenizer [8]), and the summaries are 39 tokens long.
@@
Recently, the dataset Cem Mil Podcasts was made available for non-commercial research purposes. The dataset contains transcripts and some metadata for about 114 000 podcasts in Brazilian Portuguese and another 8 000 episodes in European Portuguese. On average, the transcripts have about 9 500 words (with a median of approximately 6 700 words). The longest transcript has about 205 000 words. The task is to propose a description for the episode using the transcript as input and compare it to the one given by the creator of the podcast. To access the dataset, one must request permission from the maintainers. However, according to the official website [^4] https://podcastsdataset.byspotify.com/, as of December 2023, they no longer take requests to access it due to shifting priorities.
@@
Despite the availability of some summarization datasets in Portuguese, there is a lack of open datasets for summarizing long documents. This is a gap that PublicHearingBR fills. It includes 206 long documents and their summaries. The documents are transcripts of public hearings of the Brazilian Chamber of Deputies on various topics and the task is to extract the relevant individuals and their opinions, which can be used to write a news article about the event. The summaries are metadata indicating the individuals and what they said or supported, and a related news article about the public hearing. The dataset is open and can be freely accessed [^5] https://huggingface.co/datasets/unicamp-dl/PublicHearingBR. Besides, PublicHearingBR differs from other Portuguese datasets not only due to the length of the documents to be summarized. The generated summary is not limited to a sentence or decision: it lists the main relevant points of the debate contained in the document.
3 PublicHearingBR- A long document summarization dataset
@@ 3.1 Concept and purpose of the dataset
The Committees of the Brazilian Chamber of Deputies conduct public hearings and debates with civil society entities on various topics of national interest. The transcripts of these hearings are published online as public documents and are freely available on their official website [^6] https://www.camara.leg.br. The transcripts are extensive documents, containing dozens of pages and sometimes exceeding a hundred.
The Agência Câmara de Notícias, the news agency of the Brazilian Chamber of Deputies, often publishes news articles about these public hearings. The articles are written by journalists who follow the daily activities of the Chamber of Deputies and, after listening to the hearings, synthesize the debates into news articles. The journalists consider not only the content of the hearings but also the social and political context to focus on the most relevant points of the discussions.
The article, a short document, can be seen as a summary of a long document, the transcript. From the article, we can extract a structured summary (metadata) indicating the main topic of the hearing, the individuals participating in the debates, as well as their opinions or statements. PublicHearingBR groups this information into triples, meaning each sample contains a public hearing transcript, a news article, and metadata which link the opinions to the individuals mentioned in the document. As the article was written by a specialized team, we can consider its text and the metadata extracted from it as ground truth (gold standard).
@@
    Summarizing long documents to get the most relevant opinions of the individuals: The transcript can be used to extract the main individuals involved in the public hearing and their opinions. The extracted data can be compared with the metadata in the dataset.
@@
    Writing a journalistic article using a set of metadata: In this case, the metadata can be used as input for a pipeline to write an article, which should be compared with the article available in the dataset.
@@
    Writing a journalistic article using a long document: The transcript can be used in a pipeline to write a journalistic article, which can be compared with the article available in the dataset.
@@ 3.2 Step-by-step process for creating PublicHearingBR
Figure 1 shows the flowchart for the creation of the dataset. The process began with a manual selection of URLs of news articles written by the Agência Câmara de Notícias that were related to public hearings. During this step, we selected 206 articles about public hearings that occurred between November 2021 and May 2024. Then, we manually searched the websites of the Committees that held the hearings to find the URLs of the transcripts.
The next step was to download of HTML files for both the selected articles and the transcripts and parse them. The post-processing includes cleaning the HTML structure and isolating the main text (videos and figures of the articles were removed).
The post-processed text of the articles was used as input to a LLM to extract the structured summary (metadata). We used the OpenAI’s GPT-4o model with the prompt shown in Figure 3777All the prompts used in this article are available in Appendixes A (original version, in Portuguese) and B (translated version, in English).. The generated metadata includes the main topic of the hearing, the individuals participating in the hearings, and their opinions.
At this point, we have data containing a list of opinions expressed by each individual mentioned in the articles. To ensure the quality of the dataset, we manually reviewed all the extracted opinions and, during this analysis, we identified that some opinions in the list meant the same thing. This occurred because the news articles that we are working with in this study often describe an opinion and use a direct quotation to exemplify it. As a result, the automatic extraction considered this as two separate opinions. Therefore, we manually corrected all opinions with this characteristic, i.e., we combined these distinct sentences into a single opinion.
The final step was to merge all this information into a single JSONL file. Thus, the generated dataset contains the transcript of the public hearing (long document), the news article (summary in unstructured format), and the extracted and manually corrected metadata from the articles (summary in structured format).
@@ 3.3 Characteristics of PublicHearingBR
PublicHearingBR contains 206 samples. Each sample includes a public hearing transcript, a news article, and a dictionary containing the metadata that links opinions with individuals as presented in the article. Table 1 shows the statistics of the word counts in the transcripts and articles, and the number of individuals and opinions in the metadata.
@@
On average, the transcripts have about 18 000 words and 90% of them have between approximately 8 500 and 30 500 words. There are two transcripts with over 50 000 words, and the longest transcript available in the dataset has nearly 150 000 words.
The articles have a more uniform length, with 627 words on average. The longest article has approximately twice this length. The low variability in length is characteristic of news articles.
An average article is typically about 4% of the length of an average transcript, resulting in a compression rate of about 96%. Due to the relatively consistent length of articles and the varying length of transcripts, an article can be up to 15% of the length of a transcript.
On average, each article in the dataset presents the opinions of approximately five individuals, and there are about ten opinions per article. The articles have at most 31 opinions.
4 Experiment - Summarization of public hearing transcripts using ChatGPT

### v2

@@ Abstract
This paper introduces PublicHearingBR, a Brazilian Portuguese dataset designed for summarizing long documents. The dataset consists of transcripts of public hearings held by the Brazilian Chamber of Deputies, paired with news articles and structured summaries containing the individuals participating in the hearing and their statements or opinions. The dataset supports the development and evaluation of long document summarization systems in Portuguese. Our contributions include the dataset, a hybrid summarization system to establish a baseline for future studies, and a discussion of evaluation metrics for summarization involving large language models, addressing the challenge of hallucination in the generated summaries. As a result of this discussion, the dataset also includes annotated data to evaluate natural language inference tasks in Portuguese.
@@ 1 Introduction
With huge volumes of text being added daily to information systems, it is difficult and time-consuming for users to search through this material, which often includes a mix of relevant and irrelevant information [1, 2]. Thus, in many contexts, it is more practical to access a summary that contains the most relevant information from the data.
@@
Text summarization is a natural language processing task that reduces a text to a shorter version while retaining its main ideas. This task is important for various applications, including information retrieval, content management, and automatic news article writing, among others. Since manual summarization is time-consuming and often impractical, developing automatic systems to perform this task is necessary [1].
@@
Automatic text summarization techniques are generally divided into extractive, abstractive, and hybrid methods [3, 4]. Abstractive summarization generates new sentences that capture the essence of the original text, while extractive summarization selects and concatenates segments directly from the source. Hybrid methods combine elements of both strategies. Traditionally, most research has focused on extractive methods. However, in recent years, the emphasis has shifted to abstractive or hybrid approaches [1]. With advancements in generative artificial intelligence systems, these methods have become increasingly relevant.
@@
Documents involved in summarization tasks vary widely, ranging from short news articles to lengthy research papers and even collections of multiple documents. Each type presents distinct challenges. Summarizing short documents, such as news articles, is easier to perform and evaluate. Multi-document summarization typically depends on a prior retrieval stage focused on a specific topic within those documents. Summarizing long documents usually involves high compression rates, requiring careful assessment of the most relevant pieces of information. As the information may be dispersed throughout the text, this task can be complex.
@@
A typical pipeline for evaluating an automatic text summarization system includes several stages: pre-processing, summarization, post-processing, and evaluation [4]. The system is evaluated by comparing the generated summary with a reference summary using metrics such as ROUGE [5] and Check-Eval [6]. Regardless of the metric, publicly available standardized datasets are generally used to evaluate new methods.
@@
The availability of datasets is essential to advancing research and development in summarization methods [7], as they provide the necessary resources to train and evaluate new systems. However, most available datasets are in English, which limits the development of summarization tools for non-English speakers. Portuguese, one of the world’s most widely spoken languages [^1] https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers, illustrates this situation. Despite its wide use, there is a lack of summarization datasets in Portuguese, and existing ones usually focus on short documents like news articles, leaving a gap in resources for long document summarization.
@@
To address this gap, this paper introduces PublicHearingBR, an open dataset for summarizing long documents in Brazilian Portuguese. The documents are transcripts of public hearings held by the Brazilian Chamber of Deputies, during which individuals discuss a specific topic. Each hearing’s transcript is published online, and certain characteristics of these documents may present challenges for summarization, including: 1. They are extensive texts, often spanning dozens of pages and sometimes exceeding one hundred; 2. The opinions of a given individual may be dispersed throughout the text; 3. Some opinions of a participant may be inferred from their agreement or disagreement with others; and 4. The topics discussed are diverse, ranging from socially controversial issues (e.g., vaccination mandates, minority rights, and climate change) to highly specific topics (e.g., regulation of online platforms and installation of offshore wind farms).
@@
PublicHearingBR consists of 206 pairs of transcripts with corresponding news articles and metadata that indicate the participants in the hearing and their opinions (a structured summary). The articles were extracted from the Agência Câmara de Notícias [^2] https://www.camara.leg.br/noticias, the news agency of the Brazilian Chamber of Deputies. Thus, we consider these articles to be adequate summaries of the hearings. The structured summaries were extracted from the articles.
@@
To test the dataset, we used ChatGPT with a custom prompt [^3] https://openai.com/index/introducing-gpts/ to extract relevant information from the transcripts using a hybrid summarization approach. The goal is not to solve the problem of summarizing public hearings, but to establish a baseline using an off-the-shelf tool available to end users. To evaluate the results, we propose using large language models (LLMs) to compare the experiment’s output with the dataset. We use LLMs both to verify whether the extracted opinions are present in the dataset (similar to the Check-Eval metric [6]) and to identify hallucinations.
@@
The discussion on hallucinations also led to manually annotated data containing 4 238 opinions, excerpts of text, and a flag indicating whether the opinions can be inferred from those excerpts. In other words, PublicHearingBR can also be used as a natural language inference (NLI) dataset for Brazilian Portuguese.
The contributions of this paper are: 1. a dataset with 206 pairs of long documents and summaries in Brazilian Portuguese, designed for long document summarization; 2. an LLM-based hybrid summarization system, which provides a baseline for future studies; 3. a discussion of evaluation metrics for summarization using LLMs, including how to identify hallucinations in the generated summaries; 4. additional annotated data derived from the previous items, consisting of 4 238 opinions, text excerpts, and a manually annotated flag indicating if the opinions can be inferred from those excerpts, enabling its use for NLI tasks in Brazilian Portuguese.
@@ 2 Related Work
The availability of datasets in Portuguese is limited, especially for text summarization. Although large corpora in Brazilian Portuguese have been published recently (e.g., [8]), they are intended for other purposes, such as fine-tuning LLMs. Existing summarization datasets in Portuguese are scarce and usually focus on short texts.
@@
One of the first summarization datasets available in Portuguese is the TeMário [9], which contains 100 pairs of journalistic news articles from two newspapers, along with summaries written by a specialist. On average, the texts contain approximately 600 words, and the summaries about 200 words. The specialist was explicitly instructed to keep the summaries between 25% and 30% of the original text length. The dataset was later expanded with an additional 150 samples, with an average of 1 200 words per article [10].
@@
Other Portuguese datasets for news summarization include the CSTNews [11] and the RecognaSumm [12]. CSTNews is a manually annotated dataset with 140 short texts. All the texts together amount to about 47 000 words – that is, just over 330 words per sample. On the other hand, RecognaSumm contains about 130 000 samples and considers the summary to be the heading and subheading of the news. However, these elements often describe the main topic of the news to capture the reader’s attention, rather than serving as an effective summary of the article.
@@
News articles are, in general, short texts. One Portuguese dataset with slightly longer documents is the RulingBR [13], which contains over 10 000 rulings from the Brazilian Supreme Federal Court. Each document includes four sections: Ementa, Relatório, Voto, and Acórdão. The Ementa section is considered the reference summary of the decision. Documents average around 1 200 words and can reach up to 2 000 words. Summaries contain up to 150 words, averaging about 90 words, which corresponds to approximately 7% of the total document length. This dataset focuses on summarizing judicial texts to extract the court’s decision.
@@
There are also multilingual summarization datasets, notably WikiLingua [14] and XL-Sum [15]. WikiLingua focuses on “How To” texts, while XL-Sum targets news articles. Both provide texts in English paired with summaries written in one or more languages (up to 18 languages for WikiLingua and up to 44 for XL-Sum). In general, the texts are short. For instance, WikiLingua articles average 391 tokens, and their summaries 39 tokens, based on the mBART tokenizer [16].
@@
Recently, the Cem Mil Podcasts dataset was released for non-commercial research. It contains transcripts and some metadata for about 114 000 podcasts in Brazilian Portuguese and 8 000 episodes in European Portuguese. On average, transcripts have about 9 500 words, with a median of approximately 6 700 words. The longest transcript has about 205 000 words. The task is to generate a description for each episode using the transcript as input and compare it to the one provided by the podcast creator. Access to the dataset requires permission from the maintainers. However, according to the official website [^4] https://podcastsdataset.byspotify.com/, as of December 2023, they are no longer accepting access requests due to shifting priorities.
@@
Despite the availability of some summarization datasets in Portuguese, there is a lack of open datasets for summarizing long documents. This is a gap that PublicHearingBR fills. It includes 206 long documents and their summaries. The documents are transcripts of public hearings held by the Brazilian Chamber of Deputies on various topics, and the task is to extract the relevant individuals and their opinions. The summaries consist of metadata indicating the individuals and what they said or supported, along with a related news article about the public hearing. The dataset is open and can be freely accessed [^5] https://huggingface.co/datasets/unicamp-dl/PublicHearingBR. Besides, PublicHearingBR differs from other Portuguese datasets not only because of the length of the documents to be summarized, but also because the summaries list the main points raised during the hearings rather than being limited to a sentence or a single decision.
3 PublicHearingBR– A long document summarization dataset
@@ 3.1 Concept and purpose of the dataset
The committees of the Brazilian Chamber of Deputies conduct public hearings on various topics of national interest. The transcripts of these hearings are lengthy public documents, often spanning dozens or even a hundred pages, and are freely available on the official website [^6] https://www.camara.leg.br.
Agência Câmara de Notícias, the news agency of the Brazilian Chamber of Deputies, regularly publishes news articles about these hearings. The articles, which are short documents, can be seen as summaries of longer ones – the transcripts. From them, we can extract structured summaries (metadata) indicating the main topic of the hearing, the participants in the debate, and their opinions or statements.
PublicHearingBR groups this information into triples, each containing a public hearing transcript, a news article, and metadata linking the opinions to the individuals who participated in the hearing. As the article was written by a specialized team, we consider both its text and the metadata extracted from it as ground truth (gold standard).
@@
    Summarizing long documents to extract the most relevant opinions of the individuals: the transcript can be used to identify the main participants in the public hearing and their opinions, which can then be compared with the metadata provided in the dataset.
@@
    Writing a news article from the metadata: the metadata can serve as input for a pipeline that writes an article, which can be compared to the one provided in the dataset.
@@
    Writing a journalistic article using a long document: the transcript can be used to write a news article, which can be compared with the one available in the dataset.
The experiments presented in this paper focus on the first use case of the dataset: summarizing long documents to extract the most relevant opinions of the individuals involved. The other two use cases involve news generation and are outside the scope of this article.
@@ 3.2 Step-by-step process for creating PublicHearingBR
Figure 1 shows the flowchart for the dataset creation process. It began with the manual selection of URLs for news articles published by Agência Câmara de Notícias related to public hearings. In this step, we selected 206 articles covering hearings held between November 2021 and May 2024. Then, we manually searched the websites of the committees that conducted the hearings to find the URLs of their transcripts.
The next step was to download the HTML files for both the selected articles and the corresponding transcripts and parse them. During post-processing, the HTML structure was removed to extract just the main text.
The raw text of the articles was used to extract the structured summary. This process could have been done entirely manually. However, to speed it up, we opted to use an LLM to extract a preliminary summary, which was then manually corrected. We used OpenAI’s GPT-4o model with the prompt shown in Figure 4777All prompts used in this article are available in Appendices A (original version, in Portuguese) and B (translated version, in English).. The generated metadata includes the main topic of the hearing, the participants, and their opinions. At the end of this stage, considering all 206 news articles, a total of 3 054 opinions expressed by the hearing participants were extracted.
To ensure the quality of the dataset, we manually reviewed all extracted opinions by reading the news articles and comparing them with the generated summary. Two main types of errors were found in the automatic extraction. The first was the attribution of statements to the wrong participant. The second was the splitting of a single opinion into two, which artificially inflated the total number of opinions. This happens because, in this type of news article, it is common to present a claim followed by a direct quote that illustrates it. All these errors were manually corrected. In cases of incorrect attributions, the opinions were reassigned to the correct participant. When a single opinion was split into two, they were merged.
At the end of this process, 2 203 opinions related to the 206 dataset samples remained. This data was made available in a single JSON file [^8] PublicHearing_LDS.jsonl, available in the repository of this article..
@@ 3.3 Characteristics of PublicHearingBR
PublicHearingBR contains 206 samples. Each sample includes a public hearing transcript (long document), a news article (unstructured summary), and a dictionary with metadata that links opinions to individuals as presented in the article (structured summary). Table 1 shows statistics on word counts in the transcripts and articles, as well as the number of individuals and opinions in the metadata.
@@
On average, the transcripts have about 18 000 words, with 90% of them ranging between approximately 8 500 and 30 500 words. Two transcripts have over 50 000 words, and the longest has nearly 150 000 words.
The articles have a more uniform length, averaging 627 words, with the longest one being roughly twice as long. The low variability in length is typical of news articles of the same type.
An average article is typically about 4% of the length of an average transcript, resulting in a compression rate of about 96%. However, due to the relatively consistent length of articles and the varying length of transcripts, an article can be up to 15% of the length of a transcript.
On average, each article in the dataset presents the opinions of about five individuals, with roughly ten opinions per article. The maximum number of opinions in an article is 31.
4 Experiment – Summarization of public hearing transcripts using ChatGPT